# Constraining the AGN formation channel for detected black hole binary mergers up to z=1.5 with the Quaia catalogue

### v1

@@ 1 Introduction
Another reason why dynamically dense environments might be the hosts of the most massive mergers detected by the LVK collaboration is the fact that their escape velocities can be large enough to retain the recoil-kicked remnant of a coalescence, turning it into a potential component of a subsequent GW event. This “hierarchical merger” scenario might be a physical interpretation for the existence of stellar-mass Black Holes (sBHs) in the pair-instability mass gap (Gerosa & Berti, 2017, 2019; Yang et al., 2019; Barrera & Bartos, 2022). This discontinuity is predicted between $\approx 50{\rm M}_{\odot}$ and $\approx 120{\rm M}_{\odot}$ in the sBH mass spectrum as a consequence of the complete disruption of the core during the supernova event at the end of the life cycle of stars with very high masses, that are therefore expected to leave no remnant (Heger & Woosley, 2002; Belczynski et al., 2016). However, the astrophysical mass distribution of sBHs predicted by the LVK collaboration shows evidence for the presence of objects in the pair-instability mass gap (Abbott et al., 2023). This challenges the hypothesis that all the merging systems detected by the LVK collaboration have originated from an isolated stellar binary.
@@
Accretion discs of AGN are a unique type of potential host environment for the assembly and the merger of binaries of compact objects. The reason is that in this so-called “AGN formation channel” the binaries and their components are not expected to interact only with other stellar objects, but also with the gas the accretion disc consists of. Thanks to this interaction, the disc might for example capture compact objects that have orbits moderately inclined with respect to its plane (Ostriker, 1983; Fabj et al., 2020; Nasim et al., 2023). These disc captures have the effect of increasing the number density of compact objects in the disc, where the binary formation can be gas-assisted (Tagawa et al., 2020; DeLaurentiis et al., 2023; Rowan et al., 2023). Another process which, as far as the formation of merging binaries of compact objects is concerned, is typical of the AGN scenario is migration: the radial motion of compact objects orbiting around the central supermassive black hole. Inward migration takes place in AGN discs where the net torque exerted by the gas onto the orbiting compact object is negative (Paardekooper et al., 2010; Bellovary et al., 2016), and can increase the number density of sBHs and neutron stars in the inner part of the accretion disc, facilitating the formation of binaries.
@@
One possible way to put constraints on the fractional contribution of the AGN channel to the total merger rate of compact object binaries is through the investigation of the spatial correlation between the sky maps of the events detected by the LVK collaboration and the positions of observed AGN. This approach has been first suggested in Bartos et al. (2017), where it was estimated that 300 GW detections are needed to statistically prove, with a significance of $3\sigma$, the GW-AGN connection if half of the mergers happened in an AGN. In Corley et al. (2019) it was later found that the number of required detections decreases by a factor $\approx 3$ if it is assumed that the inclination of the binary angular momentum with respect to the line of sight is known with an uncertainty of $5^{\circ}$. In the simulated GW detections used in Bartos et al. (2017) and Corley et al. (2019) the interferometers of LIGO and Virgo have been assumed to be working at design sensitivity. This level of instrumental sensitivity is a factor $\gtrsim 2$ better with respect to the one reached by the interferometers during the first three observing runs (Cahillane & Mansell, 2022).
@@
In Veronesi et al. (2022) the analysis presented in Bartos et al. (2017) was repeated assuming a realistic distribution of the sizes of the 90 per cent Credibility Level (CL) localisation volumes (V90) of the mock GW detections. This distribution has been created by using the sensitivity curves that characterised the LIGO and Virgo interferometers during their third observing run (O3). It was found that the amount of data collected during O3 would be enough to prove with a $3\sigma$ significance that the AGN channel contributes to half of the total merger rate only if the GW events originate form a rare host population with a number density lower than $10^{-7}{\rm Mpc}^{-3}$. Integrating the the AGN luminosity function of Hopkins et al. (2007), we find that this density corresponds to objects in the local Universe with a bolometric luminosity greater than approximately $10^{46}{\rm erg\ s}^{-1}$.
@@ 2.1 GW events
In this work we use the 159 sky maps of mergers of compact object binaries detected by the interferometers of the LVK collaboration up to June 1st, 2024. In particular, all the GW events from the first three observing runs are included, with the only exclusion of GW200308_173609 and of GW200322_091133. These two events are not taken into consideration because they are poorly localised, and the corresponding value of V90 cannot be estimated from the currently available posterior samples. Moreover, we use all the events from the first half of O4 (O4a) as well as all the events from its second half (O4b) that have been detected not later than June 1st, 2024. For this currently ongoing observing run, we select the detections that have a probability of being of terrestrial origin smaller than 1 per cent.
@@
The sky maps of the 71 events of O4 are downloaded from the Gravitational-Wave Candidate Event Database  [^1] https://gracedb.ligo.org/, operated by the LIGO Scientific Collaboration. For each event, we use the most recent version of its sky map, which is either obtained from the Bilby localisation algorithm (Ashton et al., 2019) or from the Bayestar one (Singer & Price, 2016).
@@ 2.3.1 Evaluation of bolometric luminosity from ${\rm mag}_{\rm G_{{\rm RP}}}$:
The intrinsic luminosity of each AGN emitted at a rest-frame frequency of $\nu_{{\rm G}_{\rm RP}}=10^{14.588}{\rm Hz}$ 444this value has been calculated assuming a fiducial wavelength for the ${\rm G}_{\rm RP}$ band of 7750Å. is then calculated as follows:
@@ 5 Discussion and conclusion
We calculate the posterior probability function on $f_{\rm AGN}$ for different thresholds on the bolometric luminosity. We find that this function always peaks at $f_{\rm AGN}=0$. We calculate the upper limits of the 68, 90, and 95 per cent credibility intervals on $f_{\rm AGN}$. The main results of this work are summarised in Figure 5 and in Table 3. In particular we estimate that no more than the 21 per cent of the detected GW events used in our analysis originated from an un-obscured AGN with a bolometric luminosity higher than $10^{44.5}{\rm erg\ s}^{-1}$ with 95 per cent credibility. Objects brighter than such threshold consist of almost the entirety of our Quaia sample. Fainter objects are not included due to the flux limitations of Gaia. Tighter constraints are obtained when higher luminosity thresholds are considered. In particular we find that $f_{\rm AGN}$ is not greater than 11 per cent at a CL of 95 per cent when un-obscured quasars brighter than $10^{45}{\rm erg\ s}^{-1}$ are considered.
@@
The observational constraints presented in this work consist of a generalisation of the ones obtained in Veronesi et al. (2023). Here we investigate a wider range of AGN luminosities and the entirety of the redshift range reached by the interferometers of the LVK collaboration. For this reason, in order to obtain in the future even more general results one will need an all-sky AGN catalogue with a lower threshold in flux with respect to Quaia, to be able to extend the analysis to the faint-end of the un-obscured AGN population. We estimate that no more than one GW event out of five has originated in an un-obscured AGN with a bolometric luminosity higher than $10^{44.5}{\rm erg\ s}^{-1}$, but a larger fraction might still come from fainter objects, or from obscured ones.
While the results of this work demonstrate that the efficiency of the AGN channel can already be investigated with spatial-correlation analyses using the currently available data, in the next years it will be possible to put even tighter constraints using all the events that will be detected in the rest of O4 as well as the ones that will be detected during the fifth observing run of the LVK collaboration, O5. Using more data could either reduce the upper limits on $f_{\rm AGN}$, or it could cause a shifting of the value of such parameter that maximised the posterior distribution, moving it away from zero.
@@ Acknowledgements
The authors thank Elia Pizzati for the stimulating discussions regarding quasar luminosity functions and bolometric corrections. EMR acknowledges support from ERC Grant “VEGA P.", number 101002511. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. Software: Numpy (Harris et al., 2020); Matplotlib (Hunter, 2007); SciPy (Virtanen et al., 2020); Astropy (Astropy Collaboration et al., 2013, 2018); Bilby (Ashton et al., 2019); BAYESTAR (Singer & Price, 2016); Healpy (Zonca et al., 2019).

### v2

@@ 1 Introduction
Another reason why dynamically dense environments might be the hosts of the most massive mergers detected by the LVK collaboration is the fact that their escape velocities can be large enough to retain the recoil-kicked remnant of a coalescence, turning it into a potential component of a subsequent GW event. This “hierarchical merger” scenario might be a physical interpretation for the existence of stellar-mass Black Holes (sBHs) in the pair-instability mass gap (Gerosa & Berti, 2017, 2019; Yang et al., 2019a; Barrera & Bartos, 2022). This discontinuity is predicted between $\approx 50{\rm M}_{\odot}$ and $\approx 120{\rm M}_{\odot}$ in the sBH mass spectrum as a consequence of the complete disruption of the core during the supernova event at the end of the life cycle of stars with very high masses, that are therefore expected to leave no remnant (Heger & Woosley, 2002; Belczynski et al., 2016). However, the astrophysical mass distribution of sBHs predicted by the LVK collaboration shows evidence for the presence of objects in the pair-instability mass gap (Abbott et al., 2023). This challenges the hypothesis that all the merging systems detected by the LVK collaboration have originated from an isolated stellar binary (however, see also Belczynski, 2020).
@@
Accretion discs of AGN are a unique type of potential host environment for the assembly and the merger of binaries of compact objects. The reason is that in this so-called “AGN formation channel” the binaries and their components are not expected to interact only with other stellar objects, but also with the gas the accretion disc consists of. Thanks to this interaction, the disc might for example capture compact objects that have orbits moderately inclined with respect to its plane (Ostriker, 1983; Fabj et al., 2020; Nasim et al., 2023). These disc captures have the effect of increasing the number density of compact objects in the disc, where the binary formation can be gas-assisted (Bartos et al., 2017b; Tagawa et al., 2020; DeLaurentiis et al., 2023; Rowan et al., 2023). Another process which, as far as the formation of merging binaries of compact objects is concerned, is typical of the AGN scenario is migration: the radial motion of compact objects orbiting around the central supermassive black hole. Inward migration takes place in AGN discs where the net torque exerted by the gas onto the orbiting compact object is negative (Paardekooper et al., 2010; Bellovary et al., 2016), and can increase the number density of sBHs and neutron stars in the inner part of the accretion disc, facilitating the formation of binaries.
@@
One possible way to put constraints on the fractional contribution of the AGN channel to the total merger rate of compact object binaries is through the investigation of the spatial correlation between the sky maps of the events detected by the LVK collaboration and the positions of observed AGN. This approach has been first suggested in Bartos et al. (2017a), where it was estimated that 300 GW detections are needed to statistically prove, with a significance of $3\sigma$, the GW-AGN connection if half of the mergers happened in an AGN. In Corley et al. (2019) it was later found that the number of required detections decreases by a factor $\approx 3$ if it is assumed that the inclination of the binary angular momentum with respect to the line of sight is known with an uncertainty of $5^{\circ}$. In the simulated GW detections used in Bartos et al. (2017a) and Corley et al. (2019) the interferometers of LIGO and Virgo have been assumed to be working at design sensitivity. This level of instrumental sensitivity is a factor $\gtrsim 2$ better with respect to the one reached by the interferometers during the first three observing runs (Cahillane & Mansell, 2022).
@@
In Veronesi et al. (2022) the analysis presented in Bartos et al. (2017a) was repeated assuming a realistic distribution of the sizes of the 90 per cent Credibility Level (CL) localisation volumes (V90) of the mock GW detections. This distribution has been created by using the sensitivity curves that characterised the LIGO and Virgo interferometers during their third observing run (O3). It was found that the amount of data collected during O3 would be enough to prove with a $3\sigma$ significance that the AGN channel contributes to half of the total merger rate only if the GW events originate form a rare host population with a number density lower than $10^{-7}{\rm Mpc}^{-3}$. Integrating the the AGN luminosity function of Hopkins et al. (2007), we find that this density corresponds to objects in the local Universe with a bolometric luminosity greater than approximately $10^{46}{\rm erg\ s}^{-1}$.
@@ 2.1 GW events
In this work we use the 159 sky maps of mergers of compact object binaries detected by the interferometers of the LVK collaboration up to June 1st, 2024. In particular, all the GW events from the first three observing runs are included, with the exclusion of the binary neutron star merger GW170817 and of the two binary black hole mergers GW200308_173609 and GW200322_091133. While the former has not been considered our analysis since its host galaxy has been identified thanks to the direct detection of its electro-magnetic counterpart (Abbott et al., 2017)  [^1] We decide to use in our analysis only detected GW events which lack a confident host galaxy identification. However, the non-inclusion of GW170817 is not expected to have any significant impact on our final results regarding the whole population of detected mergers., the latter events have not been taken into account because they are poorly localised, and the corresponding value of V90 cannot be estimated from the currently available posterior samples. Moreover, we use all the events from the first half of O4 (O4a) as well as all the events from its second half (O4b) that have been detected not later than June 1st, 2024. For this currently ongoing observing run, we select the detections that have a probability of being of terrestrial origin smaller than 1 per cent.
@@
The sky maps of the 71 events of O4 are downloaded from the Gravitational-Wave Candidate Event Database  [^2] https://gracedb.ligo.org/, operated by the LIGO Scientific Collaboration. For each event, we use the most recent version of its sky map, which is either obtained from the Bilby localisation algorithm (Ashton et al., 2019) or from the Bayestar one (Singer & Price, 2016). We here specify as a caveat that these maps consist of early estimates, and may differ from the ones that will be published in the future Gravitational-wave Transient Catalogue (GWTC) containing all the properties of the events detected during O4.
@@ 2.3.1 Evaluation of bolometric luminosity from ${\rm mag}_{\rm G_{{\rm RP}}}$:
The intrinsic luminosity of each AGN emitted at a rest-frame frequency of $\nu_{{\rm G}_{\rm RP}}=10^{14.588}{\rm Hz}$ 555this value has been calculated assuming a fiducial wavelength for the ${\rm G}_{\rm RP}$ band of 7750Å. is then calculated as follows:
@@ 4 Results
To estimate the statistical significance of finding a GW event with a value of $\mathcal{L}_{i}\left(f_{\rm AGN}=1\right)/\mathcal{L}_{i}\left(f_{\rm AGN}=0\right)$ like the ones that come from the cross-match of S240511i with the two sub-samples of our AGN catalogue with the highest thresholds in luminosity, we perform 500 background realisations. In each realisation we cross-match the sky maps of the 159 detected mergers with AGN catalogues obtained by assigning a random value of galactic longitude to each of the objects contained in the most luminous sub-samples of the original. We do not vary the redshifts nor the galactic latitudes of the AGN for the sake of simplicity, since on these two spatial coordinates depend the number density of the catalogue and its completeness, which are parameters used during the calculation of the posterior distribution of $f_{\rm AGN}$. In 22 (25) background realisations involving the random AGN catalogues created from the sub-sample that has a luminosity threshold of $10^{46.5}{\rm erg\ s}^{-1}$ ($10^{46}{\rm erg\ s}^{-1}$) we find a GW event that has a single-event likelihood ratio higher than the one S240511i has when cross-matched with the original sub-samples. This suggests that high values of $\mathcal{L}_{i}\left(f_{\rm AGN}=1\right)/\mathcal{L}_{i}\left(f_{\rm AGN}=0\right)$ such as the ones observed in the case of S240511i have a probability of approximately 4.4 (5) per cent to arise from random chance associations.
@@
As mentioned in Section 1, the formation of binaries in dense dynamical environment such as AGN discs might explain the detection of sBHs with masses the distribution of which extend to the pair-instability mass gap region.
Figure 7: Logarithm of the ratio between the value of the single-event likelihood at $f_{\rm AGN}=1$ and the one at $f_{\rm AGN}=0$ for the mergers detected during the first three runs of the LVK collaboration, as a function of the corresponding total binary mass in the source frame. Each panel shows the results for a different bolometric luminosity threshold. In each plot, the markers are coloured as a function of the average completeness of the AGN catalogue in the region occupied by the 90 per cent CL localisation volume. The dashed horizontal lines mark where $\mathcal{L}_{i}\left(f_{\rm AGN}=1\right)=\mathcal{L}_{i}\left(f_{\rm AGN}=0\right)$.
While the posterior distributions of the intrinsic binary parameters are yet not accessible for the O4 events, for the GW events detected in the first three observing runs it is possible to check whether there is or not a correlation between binary mass and likelihood of being of AGN origin. The results of this test are shown in Figure 7, where for each event the likelihood of having happened in an AGN is plotted as a function of the total binary mass in the source frame.
Our finding is that the merging systems with the heaviest masses do not show a clear preference for the AGN origin. These events are the least informative in the framework of our spatial correlation analysis. This is because it exists a positive correlation between the total mass of the detected merging binaries and the correspondent V90 (see Figure 4 of Veronesi et al., 2024), and the constraining power of methods like the one presented in this work is mainly driven by the best localised events (Bartos et al., 2017a).
@@ 5 Discussion and conclusion
We calculate the posterior probability function on $f_{\rm AGN}$ for different thresholds on the bolometric luminosity. We find that this function always peaks at $f_{\rm AGN}=0$. We calculate the upper limits of the 68, 90, and 95 per cent credibility intervals on $f_{\rm AGN}$. The main results of this work are summarised in Figure 5 and in Table 3. In particular we estimate that no more than the 21 per cent of the detected GW events used in our analysis originated from an un-obscured AGN with a bolometric luminosity higher than $10^{44.5}{\rm erg\ s}^{-1}$ with 95 per cent credibility. Objects brighter than such threshold consist of almost the entirety of our Quaia sample. Fainter objects are not included due to the flux limitations of Gaia. Tighter constraints are obtained when higher luminosity thresholds are considered. In fact we find that $f_{\rm AGN}$ is not greater than 11 per cent at a CL of 95 per cent when un-obscured quasars brighter than $10^{45}{\rm erg\ s}^{-1}$ are considered. The dependence of the constraining power of our analysis with respect to the luminosity threshold is primarily caused by the fact that more luminous AGN are rared than dimmer ones, and spatial-correlation analyses that make use of catalogues of rare potential host environments are characterised by lower levels of noise.
From the point of view of the theoretical modeling of binary formation in AGN discs, there is a lack of general consensus regarding the relation between the accretion rate or the luminosity of the host environment and its efficiency to produce systems that merge in the LVK band. In Yang et al. (2019b), for example, it is found that the merger rate of sBHs increases, even if weakly, as a function of the AGN accretion rate. On the other hand, in Grishin et al. (2024) it is argued that AGN with a luminosity $\gtrsim 10^{45}{\rm erg\ s}^{-1}$ are expected not to develop in their disc regions in which the net torque exerted by the gas onto the compact objects is null. In these regions, usually referred to as "migration traps", an enhanced pile-up of objects might form in less luminous AGN, making them potentially more efficient in forming binaries with respect to brighter hosts (however, see also Gilbaum et al., 2024, where it is argued that, due to gap-opening mechanisms induced by the motion of the heaviest BHs in the disc, also AGN less luminous than $10^{43.5}{\rm erg\ s}^{-1}$ are not expected to develop stable migration traps.)
@@
The observational constraints presented in this work consist of a generalisation of the ones obtained in Veronesi et al. (2023). Here we investigate a wider range of AGN luminosities and the entirety of the redshift range reached by the interferometers of the LVK collaboration. For this reason, in order to obtain in the future even more general results one will need an all-sky AGN catalogue with a lower threshold in flux with respect to Quaia, to be able to extend the analysis to the faint-end of the un-obscured AGN population. We estimate that no more than one GW event out of five has originated in an un-obscured AGN with a bolometric luminosity higher than $10^{44.5}{\rm erg\ s}^{-1}$, but a larger fraction might still come from fainter objects, or from obscured ones. This has to be taken into consideration when comparing the results presented in this work with constraints obtained via other techniques, like the comparison between the observed distributions of intrinsic parameters of the merging systems and the theoretical predictions based on physical models of binary formation in AGN discs (see e.g. Gayathri et al., 2021, where it is estimated that the BH mass spectrum inferred from the observed mergers listed in GWTC-2 is compatible with a fractional contribution of the AGN channel to the total rate of 20 per cent.)
The results of this work demonstrate that spatial-correlation analyses that use currently available data have a constraining power comparable to that of other methods that investigate the efficiency of the AGN channel. However, in the next years it will be possible to put even tighter constraints using the final release of the sky maps of all the events that will be detected during the entirety of O4 as well as the ones that will be detected during the fifth observing run of the LVK collaboration, O5. Using more data could either reduce the upper limits on $f_{\rm AGN}$, or it could cause a shifting of the value of such parameter that maximised the posterior distribution, moving it away from zero.
@@ Acknowledgements
The authors thank Elia Pizzati for the stimulating discussions regarding quasar luminosity functions and bolometric corrections, and the anonymous referee, whose comments have helped us to clarify the role our results play in the investigation over the phyiscal origin of sBHs merging systems. EMR acknowledges support from ERC Grant “VEGA P.", number 101002511. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. Software: Numpy (Harris et al., 2020); Matplotlib (Hunter, 2007); SciPy (Virtanen et al., 2020); Astropy (Astropy Collaboration et al., 2013, 2018); Bilby (Ashton et al., 2019); BAYESTAR (Singer & Price, 2016); Healpy (Zonca et al., 2019).